The average age of Kevin, Joe and nick is 19. if you remove Kevin, who is the oldest, the average age of Joe and nick is 17. How old is Kevin?
Answer by srinivas.g(540)

The average age of Kevin, Joe and Nick is 19
sum of their age : sum = average *no of persons
=19*3
= 57
average age of Joe and Nick is 17
sum of ages of joe and nick = 2*17 = 34
so , age of Kevin = sum of their age -sum of ages of joe and nick
=57-34
=23 years
result is 23 years